OpenStudy (anonymous):

Evaluate. 6[5(3-9)-1]+2 ---------------- 7(8-2)+4

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (bbypanda16):

-138

OpenStudy (kendricklamar2014):

\[ 6[5(3-9)-1]+2\] \[=6((5)(−6)−1)+2\] \[=6(−30−1)+2\] \[=−186+2\] \[=−184\]

OpenStudy (kendricklamar2014):

@bbypanda16 Direct answer are not allowed. Also, your answer is wrong!

OpenStudy (bbypanda16):

Not uhhhh

OpenStudy (kendricklamar2014):

Show your steps then...

OpenStudy (bbypanda16):

I have the same problem in class. I cant explain how I got it. It just like pops in my head.

OpenStudy (kendricklamar2014):

Huh, the problem just changed...

OpenStudy (anonymous):

srry I forgot the rest of it :( mah bad

OpenStudy (anonymous):

Booyakashaaa???

OpenStudy (anonymous):

umm

OpenStudy (bbypanda16):

Damnnn

OpenStudy (bbypanda16):

I double check its -138

OpenStudy (anonymous):

positive 138

OpenStudy (kendricklamar2014):

Here is how you solve that problem ^ : http://prntscr.com/98t7lj

OpenStudy (kendricklamar2014):

This is the last part: http://prntscr.com/98t8oz

OpenStudy (kendricklamar2014):

@Mik3y100